# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 3.1 A fixed‑point operator
We adopt a similar approach to, e.g., [24]. Firs, we define a closed ball of $\mathrm{L}^{2}(\Omega)$ centred at the origin and of given radius $r>0$
(3.1) $\mathrm{W}:=\{\widehat{p}\in\mathrm{L}^{2}(\Omega)\,:\quad\|\widehat{p}\|_{0,\Omega}\leq r\}.$
Then, for a given $(\widehat{\boldsymbol{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W}$, thanks to the assumptions on the nonlinear permeability, we can infer that the form $\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}$ (cf. (2.1)) is continuous, as well as coercive over $\mathbf{V}$
(3.2a) $\displaystyle\big{|}\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}({\boldsymbol{\varphi}},{\boldsymbol{\psi}})\big{|}$ $\leq C_{\widetilde{a}}\,\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega},$
(3.2b) $\displaystyle\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}({\boldsymbol{\psi}},{\boldsymbol{\psi}})$ $\geq\kappa_{1}\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}^{2}\quad\forall\,{\boldsymbol{\varphi}},\,{\boldsymbol{\psi}}\in\mathbf{V}.$
@@
$\mathbf{R}(\widehat{p}):=\big{(}R_{1}(\widehat{p}),(R_{2}(\widehat{p}),R_{3}(\widehat{p}))\big{)}=(\boldsymbol{\sigma},(\boldsymbol{u},\boldsymbol{\rho}))\quad\forall\,\widehat{p}\in\mathrm{W},$
with $(\boldsymbol{\sigma},(\boldsymbol{u},\boldsymbol{\rho}))\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times(\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega))$ satisfying
(3.3) $\begin{array}[]{rllll}\displaystyle a(\boldsymbol{\sigma},\boldsymbol{\tau})&+\quad\displaystyle b(\boldsymbol{\tau},(\boldsymbol{u},\boldsymbol{\rho}))&=\displaystyle H(\boldsymbol{\tau})-c(\boldsymbol{\tau},\widehat{p})&\forall\,\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),\\[4.30554pt] \displaystyle b(\boldsymbol{\sigma},(\boldsymbol{v},{\boldsymbol{\eta}}))&&=\displaystyle F(\boldsymbol{v},{\boldsymbol{\eta}})&\forall\,(\boldsymbol{v},{\boldsymbol{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),\end{array}$
@@
$\mathbf{S}(\widehat{\boldsymbol{\sigma}},\widehat{p}):=\big{(}S_{1}(\widehat{\boldsymbol{\sigma}},\widehat{p}),S_{2}(\widehat{\boldsymbol{\sigma}},\widehat{p})\big{)}=({\boldsymbol{\varphi}},p)\quad\forall\,(\widehat{\boldsymbol{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W},$
where $({\boldsymbol{\varphi}},p)$ is such that
(3.4) $\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}({\boldsymbol{\varphi}},{\boldsymbol{\psi}})&+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}},p)&=\displaystyle\widetilde{H}({\boldsymbol{\psi}})&\forall\,{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),\\[4.30554pt] \displaystyle\widetilde{b}({\boldsymbol{\varphi}},q)&-\quad\widetilde{c}(p,q)&=\displaystyle G(q)+c(\widehat{\boldsymbol{\sigma}},q)&\forall\,q\in\mathrm{L}^{2}(\Omega).\end{array}$
@@
(3.5) $\mathbf{T}(\widehat{p}):=S_{2}(R_{1}(\widehat{p}),\widehat{p}),$
it is clear that $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)$ is a solution to (2.2) if and only if $p\in\mathrm{W}$ solves the fixed-point problem
(3.6) $\mathbf{T}(p)=p.$
@@ 3.2 Well‑definedness of $\mathbf{T}$
Let $\widehat{p}\in\mathrm{W}$ (cf. (3.1)). Then, there exists a unique $(\boldsymbol{\sigma},(\boldsymbol{u},\boldsymbol{\rho}))\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)$ solution to (3.3). In addition, there exist $C_{1},\,C_{2}>0$, such that
(3.7) $\begin{array}[]{cc}\|\boldsymbol{\sigma}\|_{\mathbf{div};\Omega}\leq C_{1}\,(\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{1}{c_{a}}\,\gamma\,\|\widehat{p}\|_{0,\Omega},\\[4.30554pt] \|\boldsymbol{u}\|_{0,\Omega}+\|\boldsymbol{\rho}\|_{0,\Omega}\leq C_{2}\,(\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{1}{\beta}\Big{(}1+\dfrac{1}{\mu c_{a}}\Big{)}\,\gamma\,\|\widehat{p}\|_{0,\Omega}.\end{array}$
Proof 3.2.
@@
(3.8) $C_{1}:=\Big{(}\dfrac{1}{c_{a}}+\dfrac{1}{\beta}\Big{)}\Big{(}1+\dfrac{1}{\mu c_{a}}\Big{)}{\quad\hbox{and}\quad}C_{2}:=\dfrac{1}{\beta}\Big{(}1+\dfrac{1}{\mu c_{a}}\Big{)}\Big{(}1+\dfrac{1}{\mu\beta}\Big{)};$
we omit further details.
@@
Let $(\widehat{\boldsymbol{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W}$. Then, there exists a unique $({\boldsymbol{\varphi}},p)\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ solution to (3.4). In addition, there exist $\widetilde{C}>0$ such that
(3.9) $\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p\|_{0,\Omega}\leq\widetilde{C}\,(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}).$
Proof 3.4.
The existence of a unique solution $({\boldsymbol{\varphi}},p)$ to (3.4) is straightforward given the properties of the forms $\widetilde{a}$, $\widetilde{b}$, and $\widetilde{c}$. By examining (3.2b), (2.9), and (2.8b), we can confirm that the assumptions of [26, Th. 3.4] are satisfied. In addition, ${\boldsymbol{\varphi}}$ and $p$ satisfy the following bounds
$\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}$ $\leq\Big{(}\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{C}_{1}\sqrt{\widetilde{\gamma}}\Big{)}\Big{(}2\mathop{\rm max}\limits\{\widetilde{C}_{2},\,\widetilde{C}_{3}\}\Big{)}^{1/2}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\,\|\widehat{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}),$
$\|p\|_{0,\Omega}$ $\leq\dfrac{1}{\widetilde{\beta}}\Big{(}1+\kappa_{2}r\,\big{(}\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{C}_{1}\sqrt{\widetilde{\gamma}}\big{)}\big{(}2\mathop{\rm max}\limits\{\widetilde{C}_{2},\,\widetilde{C}_{3}\}\big{)}^{1/2}\Big{)}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\,\|\widehat{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}),$
@@
(3.10) $\widetilde{C}_{1}:=\dfrac{1}{\widetilde{\beta}}\big{(}1+\dfrac{C_{\widetilde{a}}}{\kappa_{1}}\big{)},\quad\widetilde{C}_{2}:=\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{\gamma}\,\widetilde{C}_{1}^{2}{\quad\hbox{and}\quad}\widetilde{C}_{3}:=\widetilde{C}_{1}\Big{(}1+\dfrac{C_{\widetilde{a}}}{\widetilde{\beta}}+\widetilde{C}_{1}\dfrac{C_{\widetilde{a}}^{2}\,\widetilde{\gamma}}{\widetilde{\beta}^{2}}\Big{)};$
and the above implies (3.9). We leave out additional minor details.
@@
(3.11) $\widetilde{C}\big{(}1+\,\gamma\,C_{1}\big{)}\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}\big{)}+\dfrac{\widetilde{C}}{c_{a}}\,\gamma^{2}r\leq r,$
where $C_{1}$, $\widetilde{C}_{1}$ and $\gamma$ are defined in (3.8), (3.10) and (2.5), respectively. Then, for a given $\widehat{p}\in\mathrm{W}$ (cf. (3.1)), there exists a unique $p\in\mathrm{W}$ such that $\mathbf{T}(\widehat{p})=p$.
Proof 3.6.
From Lemmas 3.1 and 3.3, we ascertain that the operators $\mathbf{R}$ and $\mathbf{S}$, respectively, are well-defined, thereby ensuring the well-definition of $\mathbf{T}$. Furthermore, from (3.7) and (3.9), for each $\widehat{p}\in\mathrm{W}$, we deduce that
@@
$\leq\widetilde{C}\,(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}})+\widetilde{C}\,\gamma\,C_{1}\,(\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{\widetilde{C}}{c_{a}}\,\gamma^{2}\|\widehat{p}\|_{0,\Omega},$
this, combined with assumption (3.11), implies $\mathbf{T}(\mathrm{W})\subseteq\mathrm{W}$, which concludes the proof.
@@
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}({\boldsymbol{\varphi}},{\boldsymbol{\psi}})&+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}},p)&=\displaystyle\widetilde{H}({\boldsymbol{\psi}})&\forall\,{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),\\[4.30554pt] \displaystyle\widetilde{b}({\boldsymbol{\varphi}},q)&&=\displaystyle G(q)+c(\widehat{\boldsymbol{\sigma}},q)+\widetilde{c}(\widehat{p},q)&\forall\,q\in\mathrm{L}^{2}(\Omega).\end{array}$

### v2

@@ 3.1 A fixed‑point operator
We adopt a similar approach to, e.g., [24]. First, we define a closed ball of $\mathrm{L}^{2}(\Omega)$ centred at the origin and of given radius $r>0$
$\mathrm{W}:=\{\widehat{p}\in\mathrm{L}^{2}(\Omega)\,:\quad\|\widehat{p}\|_{0,\Omega}\leq r\}.$ (3.1)
Then, for a given $(\widehat{\bm{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W}$, thanks to the assumptions on the nonlinear permeability, we can infer that the form $\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}$ (cf. (2.1)) is continuous, as well as coercive over $\mathbf{V}$
$\displaystyle\big{|}\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}({\bm{\varphi}},{\bm{\psi}})\big{|}$ $\leq C_{\widetilde{a}}\,\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega},$ (3.2a)
$\displaystyle\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}({\bm{\psi}},{\bm{\psi}})$ $\geq\kappa_{1}\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}^{2}\quad\forall\,{\bm{\varphi}},\,{\bm{\psi}}\in\mathbf{V}.$ (3.2b)
@@
$\mathbf{R}(\widehat{p}):=\big{(}R_{1}(\widehat{p}),(R_{2}(\widehat{p}),R_{3}(\widehat{p}))\big{)}=(\bm{\sigma},({\bm{u}},\bm{\rho}))\quad\forall\,\widehat{p}\in\mathrm{W},$
with $(\bm{\sigma},({\bm{u}},\bm{\rho}))\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times(\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega))$ satisfying
$\begin{array}[]{rllll}\displaystyle a(\bm{\sigma},\bm{\tau})&+\quad\displaystyle b(\bm{\tau},({\bm{u}},\bm{\rho}))&=\displaystyle H(\bm{\tau})-c(\bm{\tau},\widehat{p})&\forall\,\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),\\[4.30554pt] \displaystyle b(\bm{\sigma},(\mathbf{v},{\bm{\eta}}))&&=\displaystyle F(\mathbf{v},{\bm{\eta}})&\forall\,(\mathbf{v},{\bm{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),\end{array}$ (3.3)
@@
$\mathbf{S}(\widehat{\bm{\sigma}},\widehat{p}):=\big{(}S_{1}(\widehat{\bm{\sigma}},\widehat{p}),S_{2}(\widehat{\bm{\sigma}},\widehat{p})\big{)}=({\bm{\varphi}},p)\quad\forall\,(\widehat{\bm{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W},$
where $({\bm{\varphi}},p)$ is such that
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}({\bm{\varphi}},{\bm{\psi}})&+\quad\displaystyle\widetilde{b}({\bm{\psi}},p)&=\displaystyle\widetilde{H}({\bm{\psi}})&\forall\,{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),\\[4.30554pt] \displaystyle\widetilde{b}({\bm{\varphi}},q)&-\quad\widetilde{c}(p,q)&=\displaystyle{\tilde{F}}(q)+c(\widehat{\bm{\sigma}},q)&\forall\,q\in\mathrm{L}^{2}(\Omega).\end{array}$ (3.4)
@@
$\mathbf{T}(\widehat{p}):=S_{2}(R_{1}(\widehat{p}),\widehat{p}),$ (3.5)
it is clear that $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)$ is a solution to (2.2) if and only if $p\in\mathrm{W}$ solves the fixed-point problem
$\mathbf{T}(p)=p.$ (3.6)
@@ 3.2 Well‑definedness of $\mathbf{T}$
Let $\widehat{p}\in\mathrm{W}$ (cf. (3.1)). Then, there exists a unique $(\bm{\sigma},({\bm{u}},\bm{\rho}))\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)$ solution to (3.3). In addition, there exist $C_{1},\,C_{2}>0$, such that
$\begin{array}[]{cc}\|\bm{\sigma}\|_{\mathbf{div};\Omega}\leq C_{1}\,(\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})+\dfrac{1}{c_{a}}\,\gamma\,\|\widehat{p}\|_{0,\Omega},\\[4.30554pt] \|{\bm{u}}\|_{0,\Omega}+\|\bm{\rho}\|_{0,\Omega}\leq C_{2}\,(\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})+\dfrac{1}{\beta}\Big{(}1+\dfrac{1}{\mu c_{a}}\Big{)}\,\gamma\,\|\widehat{p}\|_{0,\Omega}.\end{array}$ (3.7)
Proof.
@@
$C_{1}:=\Big{(}\dfrac{1}{c_{a}}+\dfrac{1}{\beta}\Big{)}\Big{(}1+\dfrac{1}{\mu c_{a}}\Big{)}{\quad\hbox{and}\quad}C_{2}:=\dfrac{1}{\beta}\Big{(}1+\dfrac{1}{\mu c_{a}}\Big{)}\Big{(}1+\dfrac{1}{\mu\beta}\Big{)};$ (3.8)
we omit further details. ∎
@@
Let $(\widehat{\bm{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W}$. Then, there exists a unique $({\bm{\varphi}},p)\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ solution to (3.4). In addition, there exist $\widetilde{C}>0$ such that
$\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p\|_{0,\Omega}\leq\widetilde{C}\,(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\bm{\sigma}}\|_{\mathbf{div};\Omega}).$ (3.9)
Proof.
The existence of a unique solution $({\bm{\varphi}},p)$ to (3.4) is straightforward given the properties of the forms $\widetilde{a}$, $\widetilde{b}$, and $\widetilde{c}$. By examining (3.2b), (2.9), and (2.8b), we can confirm that the assumptions of [26, Th. 3.4] are satisfied. In addition, ${\bm{\varphi}}$ and $p$ satisfy the following bounds
$\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}$ $\leq\Big{(}\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{C}_{1}\sqrt{\widetilde{\gamma}}\Big{)}\Big{(}2\mathop{\rm max}\limits\{\widetilde{C}_{2},\,\widetilde{C}_{3}\}\Big{)}^{1/2}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\,\|\widehat{\bm{\sigma}}\|_{\mathbf{div};\Omega}),$
$\|p\|_{0,\Omega}$ $\leq\dfrac{1}{\widetilde{\beta}}\Big{(}1+\kappa_{2}r\,\big{(}\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{C}_{1}\sqrt{\widetilde{\gamma}}\big{)}\big{(}2\mathop{\rm max}\limits\{\widetilde{C}_{2},\,\widetilde{C}_{3}\}\big{)}^{1/2}\Big{)}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\,\|\widehat{\bm{\sigma}}\|_{\mathbf{div};\Omega}),$
@@
$\widetilde{C}_{1}:=\dfrac{1}{\widetilde{\beta}}\big{(}1+\dfrac{C_{\widetilde{a}}}{\kappa_{1}}\big{)},\quad\widetilde{C}_{2}:=\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{\gamma}\,\widetilde{C}_{1}^{2}{\quad\hbox{and}\quad}\widetilde{C}_{3}:=\widetilde{C}_{1}\Big{(}1+\dfrac{C_{\widetilde{a}}}{\widetilde{\beta}}+\widetilde{C}_{1}\dfrac{C_{\widetilde{a}}^{2}\,\widetilde{\gamma}}{\widetilde{\beta}^{2}}\Big{)};$
and the above implies (3.9) with
$\widetilde{C}:=\Big{(}\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{C}_{1}\sqrt{\widetilde{\gamma}}\Big{)}\Big{(}2\mathop{\rm max}\limits\{\widetilde{C}_{2},\,\widetilde{C}_{3}\}\Big{)}^{1/2}+\dfrac{1}{\widetilde{\beta}}\Big{(}1+\kappa_{2}r\,\big{(}\dfrac{1}{\kappa_{1}}+\widetilde{C}_{1}+\widetilde{C}_{1}\sqrt{\widetilde{\gamma}}\big{)}\big{(}2\mathop{\rm max}\limits\{\widetilde{C}_{2},\,\widetilde{C}_{3}\}\big{)}^{1/2}\Big{)}.$ (3.10)
We leave out additional minor details. ∎
@@
$\widetilde{C}\big{(}1+\,\gamma\,C_{1}\big{)}\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}\big{)}+\dfrac{\widetilde{C}}{c_{a}}\,\gamma^{2}r\leq r,$ (3.11)
where $C_{1}$, $\widetilde{C}$, and $\gamma$ are defined in (3.8), (3.10) and (2.5), respectively. Then, for a given $\widehat{p}\in\mathrm{W}$ (cf. (3.1)), there exists a unique $p\in\mathrm{W}$ such that $\mathbf{T}(\widehat{p})=p$.
Proof.
From Lemmas 3.1 and 3.2, we ascertain that the operators $\mathbf{R}$ and $\mathbf{S}$, respectively, are well-defined, thereby ensuring the well-definition of $\mathbf{T}$. Furthermore, from (3.7) and (3.9), for each $\widehat{p}\in\mathrm{W}$, we deduce that
@@
$\leq\widetilde{C}\,(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}})+\widetilde{C}\,\gamma\,C_{1}\,(\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})+\dfrac{\widetilde{C}}{c_{a}}\,\gamma^{2}\|\widehat{p}\|_{0,\Omega},$
this, combined with assumption (3.11), implies $\mathbf{T}(\mathrm{W})\subseteq\mathrm{W}$, which concludes the proof. ∎
Remark 3.4.
Note that the argument used in Lemma 3.3 is as follows, for an arbitrary but fixed $r>0$, we define the ball $\mathrm{W}$ (cf. (3.1)). Then, for this fixed $r$, we assume that the data $\bm{f}$, $g$, ${\bm{u}}_{\mathrm{D}}$, $p_{\mathrm{D}}$, and $\gamma$ (cf. (2.5)) are sufficiently small to satisfy Hypothesis (3.11).
@@
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}({\bm{\varphi}},{\bm{\psi}})&+\quad\displaystyle\widetilde{b}({\bm{\psi}},p)&=\displaystyle\widetilde{H}({\bm{\psi}})&\forall\,{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),\\[4.30554pt] \displaystyle\widetilde{b}({\bm{\varphi}},q)&&=\displaystyle{\tilde{F}}(q)+c(\widehat{\bm{\sigma}},q)+\widetilde{c}(\widehat{p},q)&\forall\,q\in\mathrm{L}^{2}(\Omega).\end{array}$